If a and ß are the roots of the equation x 2 – x + 1 = 0, then a2009 + ß2009 =

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
the roots of given equation are \frac{1\pm i\sqrt{3}}{2}which can be represented as -w,-w2
hence -w2009+(-w)4018=1
Thanks and Regards
Shaik Aasif
askIITians Faculty